OpenStudy (anonymous):

What is the slope of 7x-y=5 ?

OpenStudy (mertsj):

1. Solve for y 2. The coefficient of x will be the slope.

OpenStudy (azureilai):

-y=7x+5 y=-7x-5 standard form is y = mx+b m = slope You can figure it out